### v1

# Polish Academy of Sciences

Abstract
The article presents an innovative approach to the visualisation of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of ”totems” whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualisations, based on synthetic test data as well as the vinhoverde 15-dimensional database on Portuguese wines, confirm the usefulness of our approach to the analysis of complex data structures.
Keywords: Multidimensional data, data visualisation, Chernoff faces, avatar glyphs, interactive 3D, dimensionality reduction, PCA, human-centred visualisation, dpVision
1 Introduction
Multidimensional data sets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them. One element which is struggling to keep up with this development is the human visual system (and human senses in general), naturally built for the 3D space, and lacking the ability to perceive, or even imagine, a space of a higher dimensionality.
There are two ways to cope with the problem:
  1. 1.
    reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data;
  2. 2.
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable pecularities of appearance. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this same capacity introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
Both approaches will be used in this work.
The type of multivariate data taken into consideration will be a cloud (finite set) of points in a multidimensional hyperspace.
The multidimensional datasets chosen for testing purposes are a) an artificial 4D shape chosen for easy visual interpretation, and b) the publicly available Vinhoverde set of data on Portuguese wines [3].
2 Symbols table
$K$
number of data points;
$n$
number of dimensions in the data to be visualized;
$m<n$
the number of dimensions the avatars are capable of representing by appearance;
$h\leq n$
number of data dimensions meaningful to the human user. By this we mean parameters which the user can relate to while thinking about the data set, because they are relevant to her work or her perception of datapoints as advantageous/disadvantageous:
$h_{s}\leq h$
number of those among them that were chosen for presentation as hyperspacial dimensions;
$h_{v}=h-h_{s}$
number of those among them that were chosen for presentation by avatar appearance;
$k\geq 3$
number of display dimensions the user is prepared to navigate. In our experiments, we limited this number to 4, so as not to overwhelm the user.
Figure 1: Data dimensions vs display dimensions
3 State of the art
Many approaches have been tested to make multivariate point clouds visible to the human observer. The human senses of vision and proprioception, and, to a lesser extent, hearing, have developed to deal with the natural 3D space, and our imagination is also strongly limited by it; the ubiquitous use of 2D screens and printed material further limits the resources at our disposal. Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
The parallel coordinate plot [4] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Fig. 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Fig. 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Fig. 2 (b)). A smoothed version of this construct is the Andrews plot (Fig. 2 (c)) [5].
A similar approach in a different (radial) coordinate system uses radial coordinates (Fig. 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Parallel and radial axis formats
A trellis chart (Fig. 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all (${N\choose{2}}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
Figure 3: A trellis chart
Purdom [6] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate ”dimensions” representing the most pertinent aspect of the data.
Figure 4: 4D hyperbox
A more condensed representation was introduced by B. Alpern and L. Carter [7]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. $n-1$ additional copies of each segment are used to form a mesh of ($N\choose{2}$ paralellograms, representing the coordinate planes of the hyperspace in a hyperbox (Fig. 4), displayed with the physical appearance of a mutifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualisation of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
In contrast, our approach introduces two important elements that, to our knowledge, are not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars — inspired by Chernoff faces — while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualisation, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualisation is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
4 Paradigm
Quite often, when multidimensional data is mentioned, we think of abstract sets of numbers to be analyzed by pure mathematics. However, there do exist data sets in which at least a part of the parameters is intuitively recognizable and can be directly interpreted in the context of a given subject area. One example is the publicly available dataset describing Portuguese wines, where find such parameters as acidulity, sugar content, or general assessment of quality. These features are readily understandable both to the expert and to the consumer. We can therefore divide multidimensional data into two categories.
The first group is those parameters that are intuitively understandable to the human observer. The values of parameters in this group refer to perceptible qualities of the objects in question, which makes them easier to interpret and compare. In the case of wines it’s the sensory and quality attributes mentioned above, to which the user can relate from his own experience. Displaying them individually as separate dimensions provides for a better analysis and understanding of the data structure.
The other group includes those parameters, which, however important they can be mathematically, have no direct reference to human perception. They can be chemical coefficients or technical aspects of the manufacturing process, whose meaning is difficult to perceive without additional context. In such cases the computer can be used to reduce their dimensionality to a smaller number of components, preserving the most essential information while reducing the burden on the visual channel.
This conceptual division between user-recognisable and abstract dimensions is central to our method. In most visualisation systems, all dimensions are treated as equal data fields, often without considering how understandable they are to the user. While this is sufficient in purely mathematical applications, it may create problems in real-world contexts, where users expect visualisations to reflect qualities they can interpret and relate to.
By assigning the more intuitive dimensions to the appearance of avatars, we use the natural strength of human perception — the ability to notice and compare visual traits. The less intuitive dimensions are projected into 3D space using classical techniques or reduced using PCA, allowing the user to explore general patterns and clusters without being distracted by unfamiliar technical details.
This dual approach makes the system flexible and adaptable to different data types and user needs. It also reflects a real-world situation, where some features (like sweetness or colour of a wine) are easier to explain and compare, while others (like pH level or sulphate concentration) may require expert interpretation. Our system helps to balance these aspects in a single coherent visualisation.
This approach merges the advantages of two tools of analysis: on one hand, the user is offered a clear and intuitive representation of the most meaningful data fields; on the other — the less intuitive ones are filtered by the computer, finally yielding a visualisation which is both information-rich and easy to interpret.
And this representation will be delivered to the user via two faculties of the human visual system: our capability of assessing spatial relationships in a 3D space, and our aptitude for perceiving facial expressions, although in a less quantitative way.
The data dimensions which were chosen for spatial display are transformed by two matrices and one scalar parameter:
  • $\mathbf{F}$
    filtering matrix which incorporates the following functions:
    1. 1.
      Selection of named parameters;
    2. 2.
      If necessary, transformation of unnamed parameters into a reduced space (PCA);
    3. 3.
      assignment of both groups to dimensions of observation space;
  • $\mathbf{V}$
    viewing matrix, describing the position and optics of the observer moving around in observation space. This matrix and the accompanying viewing algorithm determine which data points should actually be visible in the 3D viewing space, and at what position.
  • $s$
    slab thickness is the tolerance to which a data point must intercept the viewer hyperplane to be displayed. It can be set to infinity, in which case all data points will be projected onto the hyperplane.
5 implementation
The visualisation procedure takes as its input:
  1. 1.
    the number $n$ of dimensions of the cloud;
  2. 2.
    an $n*K$ data cloud (a matrix in which each of k data points is a column, and each dimension is a row);
  3. 3.
    a transformation matrix with $n$ columns and $m+k$ rows, $m$ being the number of visual parameters of an avatar.
  4. 4.
    the $(k+2)*n$ observer matrix. It is used to transform data from the $n$-dimensional data space into the k-dimensional observation space.
    Figure 5: Fragment of dialog box assigning parameters
    These transformations, projecting the data points both into the geometric visualization hyperspace and into the parameter space of the Chernoff faces, are computed from the user’s choices made in a dialog box (Fig. 5), assigning various display forms to data dimensions. Some data dimensions can be left unassigned, and they will be subjected to PCA analysis, to reduce their number to that of the available spacial dimensions.
    One of the two additional rows, the slice thickness is used to determine the distance between a given data point and the crossection hyperplane (points too far away from it are not visualized). The other extra row is needed for homogeneous representation of the data.
  5. 5.
    the viewing matrix, part of which is controlled by the user via mouse or keyboard commands, using a multidimensional compass rose to visualize the current orientation and shift. Another user-controlled parameter is the thickness of the visible slab (area between two parallel hyperplanes).
Figure 6: A point cloud intersected by a slab
Upon completion, the procedure inserts a cloud of avatars into the viewing space, i.e. the working space of dpVision [8, 9] (see Fig. 7).
Figure 7: Example of swarm of avatars with Chernoff-like faces
6 Test data
Given the difficulty of perceiving multivariate data, we need a test data set with artificially created n-dimensional points ($n>3$) whose projection onto any combination of 3 or 2 dimensions is known and recognizable on sight.
One well-known 4D solid is the tesseract [10], a four-dimensional hypercube. However, as a regular hypersolid, it’s not the best choice for our demonstrative purposes, as its orthogonal 3D projection or cross-section in any of its constituent coordinate planes would look exactly the same. We chose to disturb this symmetry by changing its aspect ratio from $1:1:1:1$ to $3:4:6:9$, a small-integer approximation of a chain of three golden ratios. It was displayed in dpVision, and a programmed procedure rotated it around five coordinate planes (in four dimensions, rotation occurs around a plane rather than a line).
Figure 8 shows the projection, into the 3D viewing space of dpVision, of this hypersolid in one of its rotated positions.
Figure 8: The anisotropic tesseract rotating in 4D, projected into a 3D viewing space.
7 Navigating the data
The four coordinate 3D spaces of the 4D hyperspace exhibit a topology homeomorphic to that of the tetrahedron. Each of them can be transformed into each of the other ones via a $90^{\circ}$ rotation (in 4D) around one of the six coordinate planes, just like each face of a tetrahedron can be transformed into any other one by a rotation (in 3D) of approximately $71^{\circ}$ around one of the six edges. The principle is shown in Figure 9, limited to four rotations (rather than the full six) to avoid clutter. The video file accompanying this paper shows our test set performing, in steps of $10^{\circ}$ at a time, the following drill:
Figure 9: Rotating the view in 4D.
  1. 1.
    Attribute numbers to coordinate spaces:
    1. (a)
      $XYZ$
    2. (b)
      $XYT$
    3. (c)
      $XZT$
    4. (d)
      $YZT$
  2. 2.
    make first coordinate space the viewing space; place two of its axes parallel to the edges of the screen.
  3. 3.
    in steps of $10^{\circ}$, rotate the cloud in 3D $90^{\circ}$ around the axis parallel to the vertical edges of the screen, then to the horizontal.
  4. 4.
    rotate the cloud in 4D, also in nine $10^{\circ}$ steps, around one of the coordinate planes, so that the next coordinate 3D space coincides with the viewing space
  5. 5.
    repeat from step 3 until all the coordinate spaces have been viewed.
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinhoverde wine database.
Figure 10: Wines data points
In this test Figure 10), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulphates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 Conclusion and Contribution
This paper introduces a new method for visualising high-dimensional data. It uses two natural abilities of the human visual system: understanding of 3D space and recognition of facial features. The method combines a 4D point cloud (rotating in 4D and projected onto a 3D space so as to make its structure accessible to human senses) with expressive icons inspired by Chernoff faces. In this way, it helps people to understand many-dimensional data in a more intuitive and user-friendly way. The 4D cloud can readily be expanded to higher dimensions, but we decided it might overwhelm the user.
The main idea is to divide the data dimensions into two groups. The first group contains features that are easy to understand — such as sensory qualities or user ratings — and these are shown as visual traits of the avatars. The second group includes more abstract or technical parameters, which are shown using 3D positions. If needed, methods such as Principal Component Analysis (PCA) are used to reduce the number of these less intuitive dimensions.
The method has been implemented as a plugin for the open-source platform dpVision. This plugin lets users explore a cloud of avatars in 3D space, where each avatar represents one data point. The user can move around the scene, change the direction of viewing, and adjust the thickness of the visible ”slab” of data, i.e. the extent, in the fourth dimension, of data that will be projected into the 3D viewing space. The assignment of data dimensions to position or appearance can also be changed.
The contribution of this work is threefold. First, it offers a new way to show complex data using both spatial position and visual appearance. Second, it provides an open and working software tool that can be used in practice. Third, it shows the method in action with two datasets: a synthetic cuboid-shaped 4D point cloud and a real-world wine quality dataset with 15 dimensions.
Our experiments show that this approach can help users to better understand complicated data structures. It can be especially useful when not all dimensions are equally important or easy to explain. Since the system is flexible and modular, it can be expanded in the future — for example, by using different avatar types or automatic selection of features.
This method can be helpful in many fields where people need to interpret data visually — such as medicine, quality control, or social sciences — especially when they are not experts in mathematics but still need to make informed decisions.
Our work proposes a new way of thinking about the structure of multidimensional data. Instead of treating all dimensions in the same way, we divide them into two categories: those that are easy to understand and those that are more technical or abstract. This division is then reflected in the visual representation. Dimensions that are more intuitive to the user are shown directly through the appearance of avatars — their shape, size, or facial expression. The other dimensions are either projected into 3D space or reduced using mathematical tools such as PCA.
We believe this distinction is important in many practical situations, where users want to understand what the data means, not just how it is mathematically structured. Unlike traditional visualisation methods, our system provides both spatial and visual channels of information, and assigns them to different types of data in a meaningful way.
To our knowledge, this is the first system that combines avatar-based glyphs with interactive 3D exploration in a way that reflects the human ability to interpret specific features more easily than others. The method has been successfully implemented in the dpVision platform and tested on both synthetic and real-world datasets. This shows that the approach is not only theoretically interesting, but also technically feasible and practically useful.
9 Future Work
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs — such as neutral geometric shapes or biologically inspired forms like trees — could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user.
Another area of development is the automation of dimension assignment. At present, the mapping of data dimensions to either spatial coordinates or avatar features is performed manually. In future versions, this process could be assisted or fully automated using heuristics or machine learning methods that take into account feature importance, variance, or user-defined priorities.
A further line of improvement concerns scalability. While the current implementation is effective for datasets with several hundred or a few thousand points, visual clutter becomes an issue when the number of avatars exceeds a certain threshold. Mechanisms such as dynamic filtering, clustering, or zoomable detail views could be added to improve usability for larger datasets.
Finally, a formal evaluation of the system is planned. Although the visualisation method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
Acknowledgements
The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view. The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Biography Section
Leszek Luchowski (D.Sc.)
has been with the IITiS PAN (formerly ZSAK PAN) since 1983, and part of its Vision Systems Group since 1987. He led the team for a part of this time, and was responsible for a number of research projects in the theory and applications of computer vision. He sustained his PhD thesis in image segmentation in 1994, and his DSc (higher doctorate) in 2014.
Dariusz Pojda (PhD)
specialises in programming (C++, Python, Java and others) and code optimisation. He is interested in data visualisation issues and is the main developer of the dpVision software. He has participated in several research projects on the application of computer vision in fields such as cultural heritage archiving, medicine, and forensics. He has also been involved in work on supporting manufacturing processes using AR/VR systems and applying 3D graphics to visualise quantum computing.

### v2

@@
 Abstract
-The article presents an innovative approach to the visualisation of multidimensional data, using icons inspired by Chernoff faces.
+The article presents an innovative approach to the visualization of multidimensional data, using icons inspired by Chernoff faces.
 The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions.
+We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a hyperspace of four, or potentially more dimensions.
 The technique is implemented as a plugin to the dpVision open-source image handling platform.
-The plugin allows the data to be interactively explored in the form of a swarm of ”totems” whose position in hyperspace as well as facial features represent various aspects of the data.
-Sample visualisations, based on synthetic test data as well as the vinhoverde 15-dimensional database on Portuguese wines, confirm the usefulness of our approach to the analysis of complex data structures.
-Keywords: Multidimensional data, data visualisation, Chernoff faces, avatar glyphs, interactive 3D, dimensionality reduction, PCA, human-centred visualisation, dpVision
+The plugin allows the data to be interactively explored in the form of a swarm of avatars whose position in hyperspace as well as facial features represent various aspects of the data.
+Sample visualizations, based on synthetic test data as well as the 12-dimensional database on Portuguese Vinho Verde wines, confirm the usefulness of our approach to the analysis of complex data structures.
 1 Introduction
@@
     increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates.
-Those additional properties can include color, size, shape parameters, any quantifiable pecularities of appearance.
+Those additional properties can include color, size, shape parameters, any quantifiable visual peculiarities.
 Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar.
@@
 One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions.
-However, this same capacity introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces.
+However, this capacity also introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces.
 Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis.
@@
 Trees are very familiar to humans, but seldom evoke strong emotions.
-Both approaches will be used in this work.
-The type of multivariate data taken into consideration will be a cloud (finite set) of points in a multidimensional hyperspace.
-The multidimensional datasets chosen for testing purposes are a) an artificial 4D shape chosen for easy visual interpretation, and b) the publicly available Vinhoverde set of data on Portuguese wines [3].
-2 Symbols table
-$K$
-number of data points;
-$n$
-number of dimensions in the data to be visualized;
-$m<n$
-the number of dimensions the avatars are capable of representing by appearance;
-$h\leq n$
-number of data dimensions meaningful to the human user.
-By this we mean parameters which the user can relate to while thinking about the data set, because they are relevant to her work or her perception of datapoints as advantageous/disadvantageous:
-$h_{s}\leq h$
-number of those among them that were chosen for presentation as hyperspacial dimensions;
-$h_{v}=h-h_{s}$
-number of those among them that were chosen for presentation by avatar appearance;
-$k\geq 3$
-number of display dimensions the user is prepared to navigate. In our experiments, we limited this number to 4, so as not to overwhelm the user.
+In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical.
+Users may assign meaningful dimensions to either geometric space or avatar visual features.
+If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via PCA to generate missing spatial or visual features.
+This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1).
+By merging spatial navigation with avatar-based representation, our plugin provides a human-centered and intuitive visualization of complex multidimensional data.
 Figure 1: Data dimensions vs display dimensions
+To validate the proposed approach, we use two types of multivariate datasets: a) an artificial 4D structure designed for easy geometric interpretation, and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
 2 STATE OF THE ART
@@
 Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
-The parallel coordinate plot [4] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Fig.
-2 (b, c)).
-An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Fig.
-2 (a)).
-When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Fig.
-2 (b)).
-A smoothed version of this construct is the Andrews plot (Fig.
-2 (c)) [5].
-A similar approach in a different (radial) coordinate system uses radial coordinates (Fig.
-2 (d)).
+The parallel coordinate plot [7] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Figure 2 (b, c)).
+An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Figure 2 (a)).
+When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Figure 2 (b)).
+A smoothed version of this construct is the Andrews plot (Figure 2 (c)) [8].
+A similar approach in a different (radial) coordinate system uses radial coordinates (Figure 2 (d)).
 While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
@@
 Figure 2: Parallel and radial axis formats
-A trellis chart (Fig.
-3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens.
-A classic setup is a full set of all (${N\choose{2}}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
+A trellis chart (Figure 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens.
+A classic setup is a full set of all $\binom{N}{2}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
 Figure 3: A trellis chart
-Purdom [6] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete and continuous.
-Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate ”dimensions” representing the most pertinent aspect of the data.
+Purdom [9] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete, and continuous.
+Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate "dimensions" representing the most pertinent aspect of the data.
 Figure 4: 4D hyperbox
@@
 They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees.
-$n-1$ additional copies of each segment are used to form a mesh of ($N\choose{2}$ paralellograms, representing the coordinate planes of the hyperspace in a hyperbox (Fig.
-4), displayed with the physical appearance of a mutifaceted 3D convex polyhedron in perspective.
-Despite the large number of methods proposed for the visualisation of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs.
+The $n-1$ additional copies of each segment are used to form a mesh of $\binom{N}{2}$ parallelograms, representing the coordinate planes of the hyperspace in a hyperbox (Figure 4), displayed with the physical appearance of a multifaceted 3D convex polyhedron in perspective.
+Despite the large number of methods proposed for the visualization of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs.
 These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
-In contrast, our approach introduces two important elements that, to our knowledge, are not present in previous work.
+An alternative, cognitively grounded approach was proposed by Pflughoeft et al.
+[11], who introduced data avatars — metaphorical figures designed to visually encode multidimensional data in ways that align with human intuitive perception.
+Their work, based on dual-process cognitive theory, shows that representations tailored to intuitive understanding can improve user performance in data interpretation tasks.
+We propose an approach that introduces two complementary elements not present in previous work.
 First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients).
 The intuitive features are assigned to visual characteristics of avatars—inspired by Chernoff faces—while the abstract features are reduced or projected into 3D spatial coordinates.
-Second, we combine this concept with interactive 3D visualisation, allowing users to explore a swarm of avatars from different angles and positions.
-While face-based visualisation is known from Chernoff’s work, it has mostly been used in two-dimensional form.
+Second, we combine this concept with interactive 3D visualization, allowing users to explore a swarm of avatars from different angles and positions.
+While face-based visualization is known from Chernoff’s work, it has mostly been used in two-dimensional form.
 We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
+3 NOTATION AND TERMINOLOGY
+Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters.
+The key symbols and abbreviations are summarized in Table 1, and are referenced throughout the description of the projection procedure and implementation details.
+Table 1: Symbols and notation used throughout the paper.
+Symbol Meaning Constraints
+Input data parameters
+$N$ Total number of data points
+$n$ Total number of data dimensions
+$X$ Original data matrix $X\in\mathbb{R}^{n\times N}$
+Implementation-dependent mapping
+$k$ Number of navigable display dimensions $k\geq 3$ (typically 4)
+$m$ Number of available avatar visual features $m\leq n-k$ (currently 10)
+User’s mapping
+$h_{S}$ Number of human-meaningful dimensions mapped to space $h_{S}\leq k$
+$h_{V}$ Number of human-meaningful dimensions mapped to avatars $h_{V}\leq m$
+$h_{A}$ Number of anonymous dimensions $h_{A}\leq n-(h_{S}+h_{V})$
+Projection matrices
+$F$ Filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$
+$V$ View matrix $V\in\mathbb{R}^{(k+1)\times(k+1)}$
+Derived datasets
+$X_{\text{spatial}}$ Spatial projection of $X$ $X_{\text{spatial}}\in\mathbb{R}^{k\times N}$
+$X_{\text{visual}}$ Avatar feature mapping of $X$ $X_{\text{visual}}\in\mathbb{R}^{m\times N}$
+$X_{\text{view}}$ Spatial data after applying $V$ $X_{\text{view}}\in\mathbb{R}^{k\times N}$
 4 PARADIGM
-Quite often, when multidimensional data is mentioned, we think of abstract sets of numbers to be analyzed by pure mathematics.
-However, there do exist data sets in which at least a part of the parameters is intuitively recognizable and can be directly interpreted in the context of a given subject area.
-One example is the publicly available dataset describing Portuguese wines, where find such parameters as acidulity, sugar content, or general assessment of quality.
-These features are readily understandable both to the expert and to the consumer.
-We can therefore divide multidimensional data into two categories.
-The first group is those parameters that are intuitively understandable to the human observer.
-The values of parameters in this group refer to perceptible qualities of the objects in question, which makes them easier to interpret and compare.
-In the case of wines it’s the sensory and quality attributes mentioned above, to which the user can relate from his own experience.
-Displaying them individually as separate dimensions provides for a better analysis and understanding of the data structure.
-The other group includes those parameters, which, however important they can be mathematically, have no direct reference to human perception.
-They can be chemical coefficients or technical aspects of the manufacturing process, whose meaning is difficult to perceive without additional context.
-In such cases the computer can be used to reduce their dimensionality to a smaller number of components, preserving the most essential information while reducing the burden on the visual channel.
-This conceptual division between user-recognisable and abstract dimensions is central to our method.
-In most visualisation systems, all dimensions are treated as equal data fields, often without considering how understandable they are to the user.
-While this is sufficient in purely mathematical applications, it may create problems in real-world contexts, where users expect visualisations to reflect qualities they can interpret and relate to.
-By assigning the more intuitive dimensions to the appearance of avatars, we use the natural strength of human perception — the ability to notice and compare visual traits.
-The less intuitive dimensions are projected into 3D space using classical techniques or reduced using PCA, allowing the user to explore general patterns and clusters without being distracted by unfamiliar technical details.
-This dual approach makes the system flexible and adaptable to different data types and user needs.
-It also reflects a real-world situation, where some features (like sweetness or colour of a wine) are easier to explain and compare, while others (like pH level or sulphate concentration) may require expert interpretation.
-Our system helps to balance these aspects in a single coherent visualisation.
-This approach merges the advantages of two tools of analysis: on one hand, the user is offered a clear and intuitive representation of the most meaningful data fields; on the other — the less intuitive ones are filtered by the computer, finally yielding a visualisation which is both information-rich and easy to interpret.
-And this representation will be delivered to the user via two faculties of the human visual system: our capability of assessing spatial relationships in a 3D space, and our aptitude for perceiving facial expressions, although in a less quantitative way.
-The data dimensions which were chosen for spatial display are transformed by two matrices and one scalar parameter:
-  • $\mathbf{F}$
-    filtering matrix which incorporates the following functions:
-    1. 1.
-      Selection of named parameters;
-    2. 2.
-      If necessary, transformation of unnamed parameters into a reduced space (PCA);
-    3. 3.
-      assignment of both groups to dimensions of observation space;
-  • $\mathbf{V}$
-    viewing matrix, describing the position and optics of the observer moving around in observation space.
-This matrix and the accompanying viewing algorithm determine which data points should actually be visible in the 3D viewing space, and at what position.
-  • $s$
-    slab thickness is the tolerance to which a data point must intercept the viewer hyperplane to be displayed.
-It can be set to infinity, in which case all data points will be projected onto the hyperplane.
+Multidimensional data sets often contain a mixture of parameters: some of them intuitively understandable to humans, and others more abstract or technical.
+To maximize interpretability, we propose a semantic division of data dimensions into two groups:
+  • Intuitive dimensions, such as quality ratings or sensory properties, are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
+  • Technical dimensions, such as chemical or structural parameters, are projected into spatial coordinates in a 3D observation space after optional dimensionality reduction.
+Figure 5: Conceptual diagram illustrating the separation of multidimensional data into spatial coordinates and avatar visual features through filtering and viewing transformations.
+Let the original data set be represented by a matrix $X\in\mathbb{R}^{n\times N}$, where $n$ is the number of original dimensions and $N$ is the number of data points.
+A filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ is constructed to separate and process the dimensions, where $k$ is the number of spatial dimensions and $m$ the number of visual avatar features.
+The filtered data are:
+$X_{\text{filtered}}=F\cdot X$
+Here, each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
+Spatial coordinates are obtained from a subset $F_{\text{spatial}}$, where $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, resulting in:
+$X_{\text{spatial}}=F_{\text{spatial}}\cdot X$
+Similarly, visual parameters are derived from $F_{\text{visual}}$, where $F_{\text{visual}}\in\mathbb{R}^{m\times n}$
+$X_{\text{visual}}=F_{\text{visual}}\cdot X$
+The spatial data are further transformed using a viewing matrix $V\in\mathbb{R}^{k\times n}$ (where typically $k=3$ or $4$):
+$X_{\text{view}}=V\cdot X_{\text{spatial}}$
+This dual mapping—avatars for intuitive features, spatial positioning for abstract features—provides a flexible and human-centered way to visualize high-dimensional data.
+The conceptual framework of the proposed method, including the separation of intuitive and technical dimensions and the application of the filtering and viewing matrices, is illustrated in Figure 5.
 5 IMPLEMENTATION
-The visualisation procedure takes as its input:
-  1. 1.
-    the number $n$ of dimensions of the cloud;
-  2. 2.
-    an $n*K$ data cloud (a matrix in which each of k data points is a column, and each dimension is a row);
-  3. 3.
-    a transformation matrix with $n$ columns and $m+k$ rows, $m$ being the number of visual parameters of an avatar.
-  4. 4.
-    the $(k+2)*n$ observer matrix. It is used to transform data from the $n$-dimensional data space into the k-dimensional observation space.
-    Figure 5: Fragment of dialog box assigning parameters
-    These transformations, projecting the data points both into the geometric visualization hyperspace and into the parameter space of the Chernoff faces, are computed from the user’s choices made in a dialog box (Fig.
-5), assigning various display forms to data dimensions.
-Some data dimensions can be left unassigned, and they will be subjected to PCA analysis, to reduce their number to that of the available spacial dimensions.
-    One of the two additional rows, the slice thickness is used to determine the distance between a given data point and the crossection hyperplane (points too far away from it are not visualized).
-The other extra row is needed for homogeneous representation of the data.
-  5. 5.
-    the viewing matrix, part of which is controlled by the user via mouse or keyboard commands, using a multidimensional compass rose to visualize the current orientation and shift.
-Another user-controlled parameter is the thickness of the visible slab (area between two parallel hyperplanes).
-Figure 6: A point cloud intersected by a slab
-Upon completion, the procedure inserts a cloud of avatars into the viewing space, i.e. the working space of dpVision [8, 9] (see Fig. 7).
-Figure 7: Example of swarm of avatars with Chernoff-like faces
+Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves the construction of appropriate transformation matrices, dimensionality reduction where necessary, and the assignment of selected data dimensions to spatial coordinates or avatar visual features.
+The implementation uses the Eigen library (C++), which allows efficient operations on matrices and vectors.
+The input data are stored in a matrix of size $n\times N$:
+Eigen::MatrixXd X(n, N);
+Each column corresponds to a single point in the data space $\mathbb{R}^{n}$, and each row represents one of the $n$ dimensions in that space.
+5.1 Assigning dimensions to categories
+Note: To avoid confusion, we refer to dimensions when describing input data, and to features when referring to the visualized output.
+Figure 6: Dialog window for assigning input dimensions to feature categories: spatial, visual, unnamed, or skipped. For spatial and visual categories, a specific output feature label can be selected.
+Input dimensions can be assigned to one of three feature categories: spatial, visual, or anonymous. Additionally, any dimension may be marked as skipped and excluded from further processing.
+This assignment is performed through a dedicated dialog window (Figure 6), which enables the user to associate input dimensions with specific feature categories and, when applicable, select predefined visual or spatial feature labels.
+The complete set of available labels for spatial and visual features is presented in Table 2.
+The resulting configuration is stored in index tables.
+It is not required to assign all available features — any missing spatial or visual values are automatically interpolated during subsequent processing.
+Table 2: List of features used in the implementation
+Category Feature Description
+User-defined spatial features
+spatial X X coordinate in 4D space
+spatial Y Y coordinate in 4D space
+spatial Z Z coordinate in 4D space
+spatial T T coordinate in 4D space
+User-defined avatar facial attributes
+visual Skin_C avatar’s skin color
+visual Hair_C avatar’s hair color
+visual Eye_S distance between the avatar’s eyes
+visual Nose_L avatar’s nose length
+visual Mouth_W avatar’s mouth width
+visual Smile degree of curvature of the avatar’s smile
+visual Frown level of frowning expression
+visual Hair_L avatar’s hair length
+visual Face_Elong avatar’s face proportions
+visual Iris_C avatar’s iris color
+features not defined by the user
+anonymous used for principal component analysis (PCA)
+Features selected by the user to skip
+skipped excluded from further processing
+5.2 Filtering matrix
+The first step is to create the filtering matrix, which in the implementation has the following form:
+Eigen::MatrixXd F(k + m, n);
+F.setZero();
+To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation, these are the 4 features: X, Y, Z, T), and the following $m$ rows correspond to visual features (in our case: 10 features representing details of the avatar’s appearance).
+The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
+  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, 1 is set at the intersection of the row for that feature and the column of the original input dimension.
+  • Then, if there are still unused rows in $F$, and the user has selected anonymous features ($h_{A}>0$), a PCA is performed on the input dimensions assigned to those anonymous features.
+The $(k+m)-(h_{S}+h_{V})$ strongest components are selected, and their coefficient vectors are inserted into the remaining empty rows of $F$.
+The transformed data is calculated as the product of the filtering matrix and the input matrix, followed by normalization.
+The procedure for creating the filtering matrix and transforming the input data is shown in Algorithm 1.
+Algorithm 1 Implementation of Data Projection
+0:  $X$ (all features), $Asgn$ (assignment of features to groups)
+0:  $X_{\text{filtered}}$ (projected data)
+1:  Initialize $F$ (zero matrix)
+2:  for each feature index $j$ do
+3:     if assignment of feature $j$ is ignored then
+4:        Skip
+5:     else if assignment of feature $j$ is spatial then
+6:        Mark corresponding spatial row in $F$
+7:     else if assignment of feature $j$ is visual then
+8:        Mark corresponding visual row in $F$
+9:     else if assignment of feature $j$ is anonymous then
+10:        Save index of feature $j$ for PCA
+11:     end if
+12:  end for
+13:  if number of mapped features $(h_{S}+h_{V})<(k+m)$ then
+14:     Perform SVD on anonymous features
+15:     Select $(k+m)-(h_{S}+h_{V})$ top components
+16:     Insert principal directions into empty rows in $F$
+17:  end if
+18:  $X_{\text{filtered}}=F\times X$
+19:  for each row in $X_{\text{filtered}}$ do
+20:     Center and scale (standardize)
+21:     if row corresponds to spatial dimension then
+22:        Shift by $+0.5$
+23:     end if
+24:  end for
+25:  return  $X_{\text{filtered}}$
+In this way, we obtain a table of $N$ points, each with $(k+m)$ dimensions. These dimensions can be divided into two categories:
+  • The first $k$ rows form the matrix X_spatial, which contains spatial features.
+  • The following $m$ rows form the matrix X_visual, which contains avatar visual features.
+Eigen::MatrixXd X_filtered(k + m, N);
+X_filtered = F * X;
+Eigen::MatrixXd X_spatial = X_filtered.topRows(k);
+Eigen::MatrixXd X_visual = X_filtered.bottomRows(m);
+The matrix $\mathbf{X}_{\text{spatial}}$ is now a ready-to-use $k$-dimensional dataset for further processing.
+5.3 View matrix
+The implementation supports navigation in a $k$-dimensional space (typically 4D) using affine transformations.
+These transformations are represented by a square matrix of size $(k+1)\times(k+1)$, which combines both rotation and translation.
+The core of this matrix is a $k\times k$ rotation matrix, generated by specifying the rotation plane (two coordinate axes) and the rotation angle.
+In the current plugin implementation, all 4D rotations are carried out in coordinate planes such as XY, XZ, XT, YZ, YT, or ZT.
+These correspond to classic SO(2) rotations within two-dimensional subspaces.
+However, the implementation can be extended to support full SO(4) rotations, offering a more expressive and flexible navigation model.
+The remaining part of the view matrix stores a translation vector and a homogeneous row to enable affine transformations.
+In code, the matrix can be built as follows:
+// Rotation only (k x k)
+Eigen::MatrixXd R = generateRotationMatrix(k, axis1, axis2, angle);
+// Create full (k+1 x k+1) view matrix with translation
+Eigen::MatrixXd V = Eigen::MatrixXd::Identity(k + 1, k + 1);
+V.topLeftCorner(k, k) = R;
+V.topRightCorner(k, 1) = translation_vector; // optional
+To apply the transformation, the spatial data must be converted into homogeneous coordinates by adding a row of ones:
+Eigen::MatrixXd X_hom(k + 1, N);
+X_hom.topRows(k) = X_spatial;
+X_hom.row(k) = Eigen::RowVectorXd::Ones(N);
+Eigen::MatrixXd X_view = V * X_hom;
+This allows smooth 4D navigation, including both rotation and panning of the data in the observation space.
+5.4 Slab-based Filtering
+To limit the number of visible data, a slab-based filtering mechanism is used.
+The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.
+The filtering uses a normal vector to the viewing space to compute signed distances from each point.
+Depending on whether the filtering is applied before or after transforming the data by the view matrix, different normal vectors are used:
+  • if filtering is done in the original spatial coordinates (X_spatial), the normal vector can be taken as the last row of the rotation matrix, which defines the orientation of the view;
+  • if filtering is done after applying the rotation (X_view), the normal vector can be defined in canonical coordinates as $[0,0,0,1]$, which corresponds to the last axis.
+Each point is projected onto the normal vector, and only those within the allowed threshold are retained.
+Eigen::VectorXd slab_values = normal_vector.transpose() * X_view;
+auto mask = slab_values.cwiseAbs().array() < slab_threshold;
+Eigen::MatrixXd X_visible = skip_masked(X_view, mask);
+Points that satisfy the slab condition—meaning they are sufficiently close to the viewing hyperplane—are passed to the next stage of the pipeline.
+This filtering effectively creates a thick slice through the multidimensional space (Figure 7).
+Figure 7: A cloud of points intersected by a slab, using a threshold of 1.5. Red points that lie outside the slab are discarded.
+5.5 3D projection
+The four-dimensional coordinates are then projected into 3D space using a standard perspective projection.
+Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform (Figure 8).
+To ensure interactive performance with large datasets, the avatars are deliberately simplified in geometry, allowing a large number of them to be rendered and manipulated in real time.
+Figure 8: Part of a sample swarm of avatars with Chernoff-like faces perspectively projected into 3D space.
 6 TEST DATA
@@
     repeat from step 3 until all the coordinate spaces have been viewed.
-This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinhoverde wine database.
-Figure 10: Wines data points
-In this test Figure 10), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulphates.
+This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinho Verde wine dataset.
+Figure 11: Example of visualization of the data from the Vinho Verde wine dataset.
+In this test (Figure 11), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulfates.
 Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose.
 The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
-8 Conclusion and Contribution
-This paper introduces a new method for visualising high-dimensional data.
-It uses two natural abilities of the human visual system: understanding of 3D space and recognition of facial features.
-The method combines a 4D point cloud (rotating in 4D and projected onto a 3D space so as to make its structure accessible to human senses) with expressive icons inspired by Chernoff faces.
-In this way, it helps people to understand many-dimensional data in a more intuitive and user-friendly way.
-The 4D cloud can readily be expanded to higher dimensions, but we decided it might overwhelm the user.
-The main idea is to divide the data dimensions into two groups.
-The first group contains features that are easy to understand — such as sensory qualities or user ratings — and these are shown as visual traits of the avatars.
-The second group includes more abstract or technical parameters, which are shown using 3D positions.
-If needed, methods such as Principal Component Analysis (PCA) are used to reduce the number of these less intuitive dimensions.
-The method has been implemented as a plugin for the open-source platform dpVision.
-This plugin lets users explore a cloud of avatars in 3D space, where each avatar represents one data point.
-The user can move around the scene, change the direction of viewing, and adjust the thickness of the visible ”slab” of data, i.e.
-the extent, in the fourth dimension, of data that will be projected into the 3D viewing space.
-The assignment of data dimensions to position or appearance can also be changed.
-The contribution of this work is threefold.
-First, it offers a new way to show complex data using both spatial position and visual appearance.
-Second, it provides an open and working software tool that can be used in practice.
-Third, it shows the method in action with two datasets: a synthetic cuboid-shaped 4D point cloud and a real-world wine quality dataset with 15 dimensions.
-Our experiments show that this approach can help users to better understand complicated data structures.
-It can be especially useful when not all dimensions are equally important or easy to explain.
-Since the system is flexible and modular, it can be expanded in the future — for example, by using different avatar types or automatic selection of features.
-This method can be helpful in many fields where people need to interpret data visually — such as medicine, quality control, or social sciences — especially when they are not experts in mathematics but still need to make informed decisions.
-Our work proposes a new way of thinking about the structure of multidimensional data.
-Instead of treating all dimensions in the same way, we divide them into two categories: those that are easy to understand and those that are more technical or abstract.
-This division is then reflected in the visual representation.
-Dimensions that are more intuitive to the user are shown directly through the appearance of avatars — their shape, size, or facial expression.
-The other dimensions are either projected into 3D space or reduced using mathematical tools such as PCA.
-We believe this distinction is important in many practical situations, where users want to understand what the data means, not just how it is mathematically structured.
-Unlike traditional visualisation methods, our system provides both spatial and visual channels of information, and assigns them to different types of data in a meaningful way.
-To our knowledge, this is the first system that combines avatar-based glyphs with interactive 3D exploration in a way that reflects the human ability to interpret specific features more easily than others.
-The method has been successfully implemented in the dpVision platform and tested on both synthetic and real-world datasets.
-This shows that the approach is not only theoretically interesting, but also technically feasible and practically useful.
+8 CONTRIBUTIONS
+This work provides several contributions to the field of multidimensional data visualization:
+  1. 1.
+    Semantic division of data dimensions
+    We propose a novel conceptual framework that divides the dimensions of a dataset into two categories:
+    • dimensions that are intuitively understandable to human users (e.g., quality ratings, sensory characteristics),
+    • and more abstract or technical dimensions (e.g., chemical or process parameters).
+    Unlike previous methods, which typically treat all dimensions equally, we assign intuitive dimensions to visual features of avatars (shape, facial expression, color) and project abstract dimensions into 3D spatial coordinates after possible dimensionality reduction.
+  2. 2.
+    Integration of avatar-based glyphs with interactive 3D navigation
+    While Chernoff faces have historically been used in static 2D representations, we integrate avatar glyphs into a fully interactive 3D environment.
+Users can freely explore the data cloud, rotate the viewpoint, adjust the visible slab, and examine clusters and patterns dynamically.
+  3. 3.
+    Practical implementation as an open-source plugin
+    We implemented our method as a working plugin for the open-source platform dpVision.
+The tool supports user-driven mapping of dimensions to visual properties and interactive spatial exploration, offering immediate practical applicability.
+  4. 4.
+    Demonstration on both synthetic and real-world datasets
+    We validated our approach using two datasets:
+    • a synthetic 4D anisotropic hypercube (for controlled geometric validation),
+    • and the Vinho Verde wine dataset (a 12-dimensional real-world dataset), illustrating the practical benefits of the method in an applied context.
+  5. 5.
+    Foundations for future extensibility
+    The visualization system is designed with modularity and flexibility in mind.
+Its architecture supports alternative avatar types, adjustable mappings of data dimensions, and dynamic filtering.
+While this version targets moderately sized datasets, its structure provides a basis for scaling to larger data collections, as further discussed in the Future Work section.
+9 CONCLUSION
+In this paper, we presented a novel method for visualizing high-dimensional data by combining two natural strengths of the human visual system: the understanding of 3D space and the recognition of facial traits.
+Our approach introduces an important distinction between intuitive and abstract data dimensions and maps them differently for optimal interpretability: intuitive features are assigned to avatar characteristics, while abstract features are reduced and projected into 3D space.
+We implemented the method as a plugin for the open-source dpVision platform and demonstrated its effectiveness on synthetic and real-world datasets.
+Preliminary results suggest that this combination of avatar-based glyphs and interactive 3D navigation offers an intuitive and flexible way to explore complex data structures, particularly when users are not domain experts in mathematics.
+Further development directions are outlined in the next section and include formal usability studies, exploration of alternative glyph styles to mitigate emotional bias, automated feature mapping, and scalability improvements to handle larger datasets.
+By distinguishing between different types of dimensions and tailoring their visual representation, our plugin proposes a human-centered paradigm that could benefit many fields where complex data must be understood and interpreted visually.
 10 FUTURE WORK
@@
 Therefore, alternative glyphs—such as neutral geometric shapes or biologically inspired forms like trees—could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user.
+Such alternatives could reduce the emotional bias associated with human-like faces and allow more neutral data interpretation depending on the application domain.
 Another area of development is the automation of dimension assignment.
@@
 A further line of improvement concerns scalability.
-While the current implementation is effective for datasets with several hundred or a few thousand points, visual clutter becomes an issue when the number of avatars exceeds a certain threshold.
-Mechanisms such as dynamic filtering, clustering, or zoomable detail views could be added to improve usability for larger datasets.
+While the current implementation is effective for moderately sized datasets (typically up to a few thousand points), its perceptual clarity decreases as the number of avatars grows.
+Glyph-based methods such as ours are not ideally suited for very large-scale datasets, as visual clutter and cognitive overload quickly become limiting factors.
+Future versions could integrate mechanisms such as dynamic filtering, clustering, sampling, or zoomable detail views to maintain interpretability in larger data collections.
+Hybrid solutions that combine glyphs with statistical overviews may also be explored.
 Finally, a formal evaluation of the system is planned.
-Although the visualisation method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains.
+Although the visualization method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains.
 These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
-Acknowledgements
+11 SUPPLEMENTARY MATERIAL
 The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view.
 The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.
-Declaration of competing interest
-The authors have no competing interests to declare that are relevant to the content of this article.
-Biography Section
-Leszek Luchowski (D.Sc.)
-has been with the IITiS PAN (formerly ZSAK PAN) since 1983, and part of its Vision Systems Group since 1987.
-He led the team for a part of this time, and was responsible for a number of research projects in the theory and applications of computer vision.
-He sustained his PhD thesis in image segmentation in 1994, and his DSc (higher doctorate) in 2014.
-Dariusz Pojda (PhD)
-specialises in programming (C++, Python, Java and others) and code optimisation.
-He is interested in data visualisation issues and is the main developer of the dpVision software.
-He has participated in several research projects on the application of computer vision in fields such as cultural heritage archiving, medicine, and forensics.
-He has also been involved in work on supporting manufacturing processes using AR/VR systems and applying 3D graphics to visualise quantum computing.

### v3

@@
-Polish Academy of Sciences
+Visualization of a multidimensional point cloud as a 3D swarm of avatars
 Abstract
-The article presents an innovative approach to the visualization of multidimensional data, using icons inspired by Chernoff faces.
-The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions.
-We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a hyperspace of four, or potentially more dimensions.
-The technique is implemented as a plugin to the dpVision open-source image handling platform.
-The plugin allows the data to be interactively explored in the form of a swarm of avatars whose position in hyperspace as well as facial features represent various aspects of the data.
-Sample visualizations, based on synthetic test data as well as the 12-dimensional database on Portuguese Vinho Verde wines, confirm the usefulness of our approach to the analysis of complex data structures.
+This paper proposes an innovative technique for representing multidimensional datasets using icons inspired by Chernoff faces.
+Our approach combines classical projection techniques with the explicit assignment of selected data dimensions to avatar (facial) features, leveraging the innate human ability to interpret facial traits.
+We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a four-dimensional (or higher) spatial embedding.
+The technique is implemented as a plugin for the open-source dpVision visualization platform, enabling users to interactively explore data in the form of a swarm of avatars whose spatial positions and visual features jointly encode various aspects of the dataset.
+Experimental results with synthetic test data and a 12-dimensional dataset of Portuguese Vinho Verde wines demonstrate that the proposed method enhances interpretability and facilitates the analysis of complex data structures.
 1 Introduction
 Multidimensional datasets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them.
-One element which is struggling to keep up with this development is the human visual system (and human senses in general), naturally built for the 3D space, and lacking the ability to perceive, or even imagine, a space of a higher dimensionality.
-There are two ways to cope with the problem:
-  1. 1.
-    reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data;
-  2. 2.
-    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates.
-Those additional properties can include color, size, shape parameters, any quantifiable visual peculiarities.
+One element that is struggling to keep up with this development is the human visual system (and human senses in general), which is naturally built for 3D space and cannot perceive or even imagine a space of higher dimensionality.
+There are two ways to cope with the problem: (i) reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data; (ii) increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates.
+Those additional properties can include color, size, shape parameters, and any quantifiable visual peculiarities.
 Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar.
-One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc.
-One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions.
-However, this capacity also introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces.
-Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis.
-To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees.
-Trees are very familiar to humans, but seldom evoke strong emotions.
+One well-known approach to avatar-based representation is the use of Chernoff faces [1], icons resembling a human face, where data values are converted into parameters such as the rounded or oblong shape of the face, hair color, and length, smiling or frowning curvature of lips, etc.
+Chernoff faces leverage the natural ability of the human observer to identify and interpret facial traits and expressions and have been widely used for intuitive visualization of multivariate data.
+However, Chernoff faces present several limitations, especially in the context of complex, high-dimensional, or interactive data analysis.
+The number of distinct visual features that can be reliably mapped onto a face is inherently limited, constraining the dimensionality that can be visualized.
+Most implementations are static and two-dimensional, limiting opportunities for dynamic exploration or spatial context.
+Importantly, the human perception of faces is subject to emotional and subjective biases—for example, certain expressions may be perceived as ”friendly” or ”untrustworthy” regardless of the underlying data [2].
+Furthermore, Chernoff faces typically treat all variables equally without distinguishing between dimensions that are intuitively meaningful to the user and those that are more abstract or technical.
+It can lead to loss of interpretability or visual clutter, especially as the number of variables increases.
+To overcome some of these issues, alternative glyphs have been proposed, such as trees [2], which are familiar to humans but evoke less emotion.
 In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical.
 Users may assign meaningful dimensions to either geometric space or avatar visual features.
-If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via PCA to generate missing spatial or visual features.
+If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via Principal Component Analysis (PCA) to generate missing spatial or visual features.
 This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1).
@@
 Figure 1: Data dimensions vs display dimensions
-To validate the proposed approach, we use two types of multivariate datasets: a) an artificial 4D structure designed for easy geometric interpretation, and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
-2 STATE OF THE ART
-Many approaches have been tested to make multivariate point clouds visible to the human observer.
-The human senses of vision and proprioception, and, to a lesser extent, hearing, have developed to deal with the natural 3D space, and our imagination is also strongly limited by it; the ubiquitous use of 2D screens and printed material further limits the resources at our disposal.
-Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
-The parallel coordinate plot [7] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Figure 2 (b, c)).
-An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Figure 2 (a)).
-When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Figure 2 (b)).
-A smoothed version of this construct is the Andrews plot (Figure 2 (c)) [8].
-A similar approach in a different (radial) coordinate system uses radial coordinates (Figure 2 (d)).
-While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
-Refer to caption Refer to caption
-(a) (b)
-Refer to caption Refer to caption
-(c) (d)
-Figure 2: Parallel and radial axis formats
-A trellis chart (Figure 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens.
-A classic setup is a full set of all $\binom{N}{2}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
-Figure 3: A trellis chart
-Purdom [9] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete, and continuous.
-Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate "dimensions" representing the most pertinent aspect of the data.
-Figure 4: 4D hyperbox
-A more condensed representation was introduced by B.
-Alpern and L.
-Carter [10].
-They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees.
-The $n-1$ additional copies of each segment are used to form a mesh of $\binom{N}{2}$ parallelograms, representing the coordinate planes of the hyperspace in a hyperbox (Figure 4), displayed with the physical appearance of a multifaceted 3D convex polyhedron in perspective.
-Despite the large number of methods proposed for the visualization of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs.
-These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
-An alternative, cognitively grounded approach was proposed by Pflughoeft et al.
-[11], who introduced data avatars — metaphorical figures designed to visually encode multidimensional data in ways that align with human intuitive perception.
-Their work, based on dual-process cognitive theory, shows that representations tailored to intuitive understanding can improve user performance in data interpretation tasks.
-We propose an approach that introduces two complementary elements not present in previous work.
-First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients).
-The intuitive features are assigned to visual characteristics of avatars—inspired by Chernoff faces—while the abstract features are reduced or projected into 3D spatial coordinates.
-Second, we combine this concept with interactive 3D visualization, allowing users to explore a swarm of avatars from different angles and positions.
-While face-based visualization is known from Chernoff’s work, it has mostly been used in two-dimensional form.
-We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
-3 NOTATION AND TERMINOLOGY
+To validate the proposed approach, we use two types of multivariate datasets: a) two synthetic datasets designed to illustrate interpretability and the effects of variable assignment—one modeling fictional politicians and another simulating a diverse set of soft drinks; and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
+The remainder of this paper is organized as follows.
+Section 2 reviews existing methods for multidimensional data visualization and highlights their limitations.
+Section 3 introduces the conceptual framework underlying our approach, including the semantic division of data dimensions, and describes the technical implementation of the visualization system as a plugin for the dpVision platform.
+Section 4 presents experiments on both synthetic and real-world datasets and explains the process of interactive exploration of the data cloud.
+Section 5 summarizes the main contributions of the work and outlines directions for future research.
+Finally, Section 6 concludes the paper.
+2 Related work
+The problem of visualizing multidimensional data has inspired a diverse set of techniques, each with its strengths and limitations.
+Projection-based methods, such as Principal Component Analysis (PCA) [7], t-distributed stochastic neighbor embedding (t-SNE) [8], and Uniform Manifold Approximation and Projection (UMAP) [9], are commonly used to reduce high-dimensional datasets to two or three dimensions, facilitating the discovery of patterns, clusters, or outliers.
+Coordinate-based visualizations, such as parallel coordinates [10] and radial plots [11], allow each variable to be explicitly represented as an axis.
+Parallel coordinates, in particular, have been widely used in medicine (e.g., ”liver profiles”), and their variants, such as Andrews plots [12], provide additional smoothing and compactness.
+These approaches are effective for moderate numbers of dimensions but can become visually cluttered as dimensionality increases.
+Matrix-based and small multiples techniques, including trellis charts [13], offer compact overviews of pairwise relationships between variables, enabling users to examine a large number of possible interactions.
+Glyph-based approaches, most notably Chernoff faces [1], map multiple variables onto visual features of an icon or avatar, leveraging innate human abilities for face and pattern recognition.
+However, the subjectivity and emotional connotations of human-like glyphs can bias interpretation, and most glyph-based methods are static and limited to two dimensions.
+Dynamic and interactive exploration techniques have also been developed, such as the grand tour and guided tour [14] and the slice tour [15].
+These methods present users with sequences of projections or thin slices through the data space, sometimes augmented with glyphs for detailed inspection (e.g., the tourr package for R).
+While powerful for uncovering high-dimensional structures, these approaches typically treat all variables as equivalent, making them challenging for non-experts to interpret and understand.
+More recently, cognitively grounded approaches have been explored.
+Pflughoeft et al.
+[16] introduced ”data avatars”—metaphorical glyphs designed to visually encode multidimensional data in ways that align with intuitive human perception.
+Their work, based on dual-process cognitive theory, demonstrated that such designs could improve interpretation; however, they did not address interactive navigation or explicit semantic division of variables.
+Compared to the above methods, our approach introduces two key innovations.
+First, we enable a semantic split of the dataset into two groups: dimensions that are intuitively meaningful to users (such as sensory or quality ratings) and those that are more technical or abstract (such as chemical or process parameters).
+The intuitive features are mapped to avatar visual traits (inspired by Chernoff faces), while the technical features are projected into spatial coordinates, optionally using dimensionality reduction (e.g., PCA).
+Second, we integrate these glyphs into an interactive 3D environment, allowing users to explore a ”swarm” of avatars, dynamically rotate the viewpoint, and apply slab-based filtering to reveal local structure.
+Our slab filtering is conceptually related to the slice tour [15], as both select points near a hyperplane for visualization.
+However, while the slice tour typically focuses on very thin slices (showing only points lying almost exactly on a hyperplane), our approach uses a slab of configurable thickness, retaining all points within a certain distance from the projection plane.
+This design supports the practical exploration of moderately sized datasets, reducing visual clutter while still revealing structural patterns.
+To our knowledge, this is the first system to combine avatar-based glyphs with interactive 3D navigation, explicit semantic variable assignment, and flexible local filtering in this manner.
+3 Methodology
 Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters.
-The key symbols and abbreviations are summarized in Table 1, and are referenced throughout the description of the projection procedure and implementation details.
+The key symbols and abbreviations are summarized in Table 1 and are referenced throughout the description of the projection procedure and implementation details.
 Table 1: Symbols and notation used throughout the paper.
@@
 $F$ Filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$
-$V$ View matrix $V\in\mathbb{R}^{(k+1)\times(k+1)}$
+$V$ View matrix: for spatial rotation in $k$-dimensional space ($k\times k$), or full affine transformation with translation in homogeneous coordinates ($(k+1)\times(k+1)$) $V\in\mathbb{R}^{k\times k}$ or $V\in\mathbb{R}^{(k+1)\times(k+1)}$
 Derived datasets
@@
 $X_{\text{view}}$ Spatial data after applying $V$ $X_{\text{view}}\in\mathbb{R}^{k\times N}$
-4 PARADIGM
-Multidimensional data sets often contain a mixture of parameters: some of them intuitively understandable to humans, and others more abstract or technical.
-To maximize interpretability, we propose a semantic division of data dimensions into two groups:
-  • Intuitive dimensions, such as quality ratings or sensory properties, are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
-  • Technical dimensions, such as chemical or structural parameters, are projected into spatial coordinates in a 3D observation space after optional dimensionality reduction.
+3.1 Paradigm and semantic assignment of variables
+In practical applications, multidimensional datasets often contain both variables that are intuitively meaningful to users (such as sensory ratings and categorical or interpretable parameters) and features that are more technical or abstract (such as chemical coefficients or structural process data).
+To maximize interpretability, our approach introduces a semantic division of data dimensions into two groups:
+  • Intuitive dimensions—readily interpreted by users—are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
+  • Technical dimensions—more abstract or difficult to interpret—are projected into spatial coordinates in the observation space, possibly after dimensionality reduction (e.g., PCA).
+The assignment of variables to these groups should be guided by domain knowledge and the analysis’s objectives.
+Features that are critical for individual profiling or human interpretation should be mapped to avatar features, while those essential for discovering structures or clusters may be included in spatial coordinates.
+Redundancy should be avoided (e.g., highly correlated features mapped to multiple avatar traits), and important variables should not be left out of both avatar and spatial assignment, as this could diminish interpretability.
 Figure 2: Conceptual diagram illustrating the separation of multidimensional data into spatial coordinates and avatar visual features through filtering and viewing transformations.
 Let the original dataset be represented by a matrix $X\in\mathbb{R}^{n\times N}$, where $n$ is the number of original dimensions and $N$ is the number of data points.
-A filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ is constructed to separate and process the dimensions, where $k$ is the number of spatial dimensions and $m$ the number of visual avatar features.
-The filtered data are:
-$X_{\text{filtered}}=F\cdot X$
-Here, each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
-Spatial coordinates are obtained from a subset $F_{\text{spatial}}$, where $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, resulting in:
-$X_{\text{spatial}}=F_{\text{spatial}}\cdot X$
-Similarly, visual parameters are derived from $F_{\text{visual}}$, where $F_{\text{visual}}\in\mathbb{R}^{m\times n}$
-$X_{\text{visual}}=F_{\text{visual}}\cdot X$
-The spatial data are further transformed using a viewing matrix $V\in\mathbb{R}^{k\times n}$ (where typically $k=3$ or $4$):
-$X_{\text{view}}=V\cdot X_{\text{spatial}}$
-This dual mapping—avatars for intuitive features, spatial positioning for abstract features—provides a flexible and human-centered way to visualize high-dimensional data.
-The conceptual framework of the proposed method, including the separation of intuitive and technical dimensions and the application of the filtering and viewing matrices, is illustrated in Figure 5.
-5 IMPLEMENTATION
-Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves the construction of appropriate transformation matrices, dimensionality reduction where necessary, and the assignment of selected data dimensions to spatial coordinates or avatar visual features.
+To implement the semantic assignment of variables, we construct a filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ that separates and processes the dimensions, where $k$ is the number of spatial dimensions and $m$ is the number of avatar visual features.
+The filtered data are given by Eq.
+(1):
+$X_{\text{filtered}}=F\cdot X.$ (1)
+Each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
+Spatial coordinates are obtained by selecting the appropriate subset $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, as shown in Eq. (2):
+$X_{\text{spatial}}=F_{\text{spatial}}\cdot X,$ (2)
+while visual parameters are derived from $F_{\text{visual}}\in\mathbb{R}^{m\times n}$, according to Eq. (3):
+$X_{\text{visual}}=F_{\text{visual}}\cdot X.$ (3)
+Finally, the spatial data are further transformed for display using a view matrix $V$, as in Eq. (4):
+$X_{\text{view}}=V\cdot X_{\text{spatial}}$ (4)
+where $V$ is typically a $k\times k$ rotation matrix.
+In practical applications, particularly for interactive 3D or 4D exploration, we use an affine $(k+1)\times(k+1)$ matrix acting on homogeneous coordinates, combining both rotation and translation (see Implementation).
+This dual mapping—avatars for intuitive features, spatial positioning for technical features—provides a flexible and human-centered way to visualize high-dimensional data (see Figure 2).
+Guidelines and limitations.
+The semantic assignment should be made carefully, considering both user needs and analytical goals.
+Important features should not be omitted or relegated solely to PCA-based projection, as their interpretability may be lost (principal components represent blends of original variables).
+The division between intuitive and technical features is context-dependent and may vary by task or user group.
+We recommend iterative testing of assignments, consultation with domain experts, and documenting mapping choices.
+When PCA is used to process unassigned (anonymous) variables, users should inspect principal component loadings to understand which original features contribute most strongly to the resulting spatial dimensions.
+Although our current tool does not include interactive visualization of these loadings, such analysis can be performed using standard statistical software.
+This framework strikes a balance between the interpretability of human-centered visualization and the structural insight provided by dimensionality reduction, remaining transparent about the trade-offs involved.
+3.2 Implementation
+Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves constructing appropriate transformation matrices, performing dimensionality reduction where necessary, and assigning selected data dimensions to spatial coordinates or avatar visual features.
 The implementation uses the Eigen library (C++), which allows efficient operations on matrices and vectors.
@@
 Each column corresponds to a single point in the data space $\mathbb{R}^{n}$, and each row represents one of the $n$ dimensions in that space.
-5.1 Assigning dimensions to categories
-Note: To avoid confusion, we refer to dimensions when describing input data, and to features when referring to the visualized output.
+3.2.1 Assigning dimensions to categories
+Note: To avoid confusion, we refer to dimensions when describing input data and to features when referring to the visualized output.
 Figure 3: Dialog window for assigning input dimensions to feature categories: spatial, visual, unnamed, or skipped. For spatial and visual categories, a specific output feature label can be selected.
@@
 skipped excluded from further processing
-5.2 Filtering matrix
-The first step is to create the filtering matrix, which in the implementation has the following form:
+3.2.2 Filtering matrix
+The first step, corresponding to Eq. (1), is to create the filtering matrix, which in the implementation has the following form:
 Eigen::MatrixXd F(k + m, n);
 F.setZero();
-To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation, these are the 4 features: X, Y, Z, T), and the following $m$ rows correspond to visual features (in our case: 10 features representing details of the avatar’s appearance).
+To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation: X, Y, Z, T), and the following $m$ rows correspond to visual features (e.g., avatar appearance details).
+This separation is formally described by Eqs.
+(2) and (3).
 The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
-  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, 1 is set at the intersection of the row for that feature and the column of the original input dimension.
+  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, one is set at the intersection of the row for that feature and the column of the original input dimension.
   • Then, if there are still unused rows in $F$, and the user has selected anonymous features ($h_{A}>0$), a PCA is performed on the input dimensions assigned to those anonymous features.
@@
 The matrix $\mathbf{X}_{\text{spatial}}$ is now a ready-to-use $k$-dimensional dataset for further processing.
-5.3 View matrix
-The implementation supports navigation in a $k$-dimensional space (typically 4D) using affine transformations.
-These transformations are represented by a square matrix of size $(k+1)\times(k+1)$, which combines both rotation and translation.
-The core of this matrix is a $k\times k$ rotation matrix, generated by specifying the rotation plane (two coordinate axes) and the rotation angle.
+3.2.3 View matrix
+The view matrix $V$ is implemented as an affine $(k+1)\times(k+1)$ matrix, combining spatial rotation (the $k\times k$ upper-left block) and translation (the final column) for interactive navigation in homogeneous coordinates.
+This approach enables efficient and flexible real-time manipulation of the data cloud in 3D or 4D space.
 In the current plugin implementation, all 4D rotations are carried out in coordinate planes such as XY, XZ, XT, YZ, YT, or ZT.
@@
 Eigen::MatrixXd X_view = V * X_hom;
-This allows smooth 4D navigation, including both rotation and panning of the data in the observation space.
-5.4 Slab-based Filtering
-To limit the number of visible data, a slab-based filtering mechanism is used.
+This corresponds directly to the transformation defined in Eq.
+(4), where the spatial coordinates are mapped into the observation space for visualization.
+This procedure enables smooth 4D navigation, including both rotation and panning of the data cloud in the observation space.
+3.2.4 Slab-based Filtering
+To limit the number of visible data, a slab-based filtering mechanism is employed.
 The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.
@@
 Figure 4: A cloud of points intersected by a slab, using a threshold of 1.5. Red points that lie outside the slab are discarded.
-5.5 3D projection
+3.2.5 3D projection
 The four-dimensional coordinates are then projected into 3D space using a standard perspective projection.
-Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform (Figure 8).
-To ensure interactive performance with large datasets, the avatars are deliberately simplified in geometry, allowing a large number of them to be rendered and manipulated in real time.
-Figure 8: Part of a sample swarm of avatars with Chernoff-like faces perspectively projected into 3D space.
-6 TEST DATA
-Given the difficulty of perceiving multivariate data, we need a test data set with artificially created n-dimensional points ($n>3$) whose projection onto any combination of 3 or 2 dimensions is known and recognizable on sight.
-One well-known 4D solid is the tesseract [12], a four-dimensional hypercube.
-However, as a regular hypersolid, it’s not the best choice for our demonstrative purposes, as its orthogonal 3D projection or cross-section in any of its constituent coordinate planes would look exactly the same.
-We chose to disturb this symmetry by changing its aspect ratio from $1:1:1:1$ to $3:4:6:9$, a small-integer approximation of a chain of three golden ratios.
-It was displayed in dpVision, and a programmed procedure rotated it around five coordinate planes (in four dimensions, rotation occurs around a plane rather than a line).
-Figure 9 shows the projection, into the 3D viewing space of dpVision, of this hypersolid in one of its rotated positions.
-Figure 9: The anisotropic tesseract rotating in 4D, projected into a 3D viewing space.
-7 NAVIGATING THE DATA
-The four coordinate 3D spaces of the 4D hyperspace exhibit a topology homeomorphic to that of the tetrahedron.
-Each of them can be transformed into each of the other ones via a $90^{\circ}$ rotation (in 4D) around one of the six coordinate planes, just like each face of a tetrahedron can be transformed into any other one by a rotation (in 3D) of approximately $71^{\circ}$ around one of the six edges.
-The principle is shown in Figure 10, limited to four rotations (rather than the full six) to avoid clutter.
-The video file accompanying this paper shows our test set performing, in steps of $10^{\circ}$ at a time, the following drill:
-Figure 10: Rotating the view in 4D.
-  1. 1.
-    Attribute numbers to coordinate spaces:
-    1. (a)
-      $XYZ$
-    2. (b)
-      $XYT$
-    3. (c)
-      $XZT$
-    4. (d)
-      $YZT$
-  2. 2.
-    make first coordinate space the viewing space; place two of its axes parallel to the edges of the screen.
-  3. 3.
-    in steps of $10^{\circ}$, rotate the cloud in 3D $90^{\circ}$ around the axis parallel to the vertical edges of the screen, then to the horizontal.
-  4. 4.
-    rotate the cloud in 4D, also in nine $10^{\circ}$ steps, around one of the coordinate planes, so that the next coordinate 3D space coincides with the viewing space
-  5. 5.
-    repeat from step 3 until all the coordinate spaces have been viewed.
-This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinho Verde wine dataset.
-Figure 11: Example of visualization of the data from the Vinho Verde wine dataset.
-In this test (Figure 11), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulfates.
-Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose.
-The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
-8 CONTRIBUTIONS
-This work provides several contributions to the field of multidimensional data visualization:
-  1. 1.
-    Semantic division of data dimensions
-    We propose a novel conceptual framework that divides the dimensions of a dataset into two categories:
-    • dimensions that are intuitively understandable to human users (e.g., quality ratings, sensory characteristics),
-    • and more abstract or technical dimensions (e.g., chemical or process parameters).
-    Unlike previous methods, which typically treat all dimensions equally, we assign intuitive dimensions to visual features of avatars (shape, facial expression, color) and project abstract dimensions into 3D spatial coordinates after possible dimensionality reduction.
-  2. 2.
-    Integration of avatar-based glyphs with interactive 3D navigation
-    While Chernoff faces have historically been used in static 2D representations, we integrate avatar glyphs into a fully interactive 3D environment.
-Users can freely explore the data cloud, rotate the viewpoint, adjust the visible slab, and examine clusters and patterns dynamically.
-  3. 3.
-    Practical implementation as an open-source plugin
-    We implemented our method as a working plugin for the open-source platform dpVision.
-The tool supports user-driven mapping of dimensions to visual properties and interactive spatial exploration, offering immediate practical applicability.
-  4. 4.
-    Demonstration on both synthetic and real-world datasets
-    We validated our approach using two datasets:
-    • a synthetic 4D anisotropic hypercube (for controlled geometric validation),
-    • and the Vinho Verde wine dataset (a 12-dimensional real-world dataset), illustrating the practical benefits of the method in an applied context.
-  5. 5.
-    Foundations for future extensibility
-    The visualization system is designed with modularity and flexibility in mind.
-Its architecture supports alternative avatar types, adjustable mappings of data dimensions, and dynamic filtering.
-While this version targets moderately sized datasets, its structure provides a basis for scaling to larger data collections, as further discussed in the Future Work section.
+Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform.
+To ensure interactive performance with large datasets, the avatars are deliberately simplified in terms of geometry, allowing a large number of them to be rendered and manipulated in real time.
+4 Results
+In this section, we present experimental results on both synthetic and real-world datasets to illustrate the interpretability and practical utility of the proposed avatar-based multidimensional visualization method.
+The analyses focus on (i) how intuitive and technical variables are assigned and visualized, (ii) what insights spatial avatars provide compared to classical approaches, and (iii) the impact of variable assignment choices.
+4.1 Small-scale illustration: Synthetic politicians dataset (N=12)
+This illustrative example utilizes a manually constructed synthetic dataset of 12 fictional politicians, each described by nine features that reflect both voter perception and performance metrics.
+Intuitive features (number of promises, fulfillment rate, public sympathy, general popularity) were mapped to avatar visual traits (e.g., smile, nose length, skin color).
+In contrast, technical features (sympathy, economic views, social views) were assigned to spatial coordinates (X, Y, Z axes).
+The remaining variables (media activity, voting effectiveness, age) were left anonymous for PCA.
+The variable group_numeric encodes four archetypes (populists, conservatives, liberals, technocrats), used only for color coding of avatar hair; each sample is labeled by name.
+Figure 5: Comparison of classical visualization methods for the synthetic politicians dataset.
+Top left: PCA scatter plot (first two components); Top right: t-SNE projection; Bottom left: parallel coordinates plot; Bottom right: Chernoff faces grid (4$\times$3) for all twelve politicians, colored and labeled by group.
+Each approach reveals different aspects of the data: scatter plots show grouping structure, parallel coordinates display multivariate profiles, and Chernoff faces summarize intuitive features—but none allows simultaneous interpretation of both structure and profile as directly as spatial avatars (see Fig.
+6).
+To provide a fair comparison with established visualization techniques, Figure 5 presents the same dataset using four classical methods: PCA scatter plot, t-SNE projection, parallel coordinates, and a 4$\times$3 grid of Chernoff faces.
+While PCA and t-SNE reveal the grouping structure, and parallel coordinates allow for the comparison of individual profiles, none of these methods offers the same level of simultaneous, intuitive insight into both position and feature profiles as spatial avatars.
+In particular, Chernoff faces communicate intuitive attributes but lack spatial context; conversely, scatter plots provide structure but obscure the source of differences between cases.
+For the twelve synthetic politicians, classical methods highlight some group separation (notably in PCA and t-SNE), but interpreting individual differences remains challenging.
+The Chernoff faces grid allows for visual comparison of intuitive features, yet it does not convey the geometric relationships between politicians or facilitate the detection of outliers in attribute space.
+This contrast underscores the unique advantage of spatial avatars in combining the interpretability of individual profiles with structural awareness.
+Variable assignment:
+Avatar features: number of promises (skin color), promise fulfillment (nose length), general popularity (smile), group affiliation (hair color).
+Spatial coordinates: sympathy (X), economic views (Y), social views (Z).
+Anonymous/PCA: media activity, voting effectiveness, age.
+Figure 6: Spatial avatar visualization of the synthetic politicians dataset.
+Avatar features encode intuitive traits (skin color: number of promises; nose length: promise fulfillment; smile: general popularity; hair color: group).
+Spatial coordinates reflect sympathy (X), economic views (Y), and social views (Z).
+Names and group colors aid interpretation.
+This visualization enables direct comparison of political archetypes, identification of outliers, and exploration of individual profiles.
+Interpretation and interactive exploration:
+The spatial avatar visualization enables immediate and intuitive comparison of both the position and the profile of each fictional politician.
+Clusters and polarization become visible in space, while characteristic features (such as promise fulfillment or popularity) are instantly ”readable” from the avatar’s face.
+Group affiliation is recognized by hair color.
+Importantly, the tool supports interactive exploration, allowing users to rotate, zoom, and navigate the scene to examine clusters, compare individuals, and discover patterns from multiple perspectives.
+This interactivity facilitates flexible hypothesis testing and nuanced understanding—features not available in static PCA scatter plots.
+PCA loadings and spatial interpretation:
+Principal Component Analysis was performed under two scenarios: (top) only on anonymous features (not mapped to avatars or spatial coordinates); (bottom) on all features assigned to spatial coordinates or left anonymous (all except avatar features).
+Figure 7 presents heatmaps of PCA loadings and the corresponding scatter plots.
+Figure 7: Effect of variable assignment on principal component analysis for the synthetic politicians dataset.
+(top) PCA on anonymous features only.
+(bottom) PCA on all features assigned to spatial coordinates or left anonymous (all except avatar features).
+Variable abbreviations: MA—media activity, VE—voting effectiveness, AG—age, SY—sympathy, EV—economic views, SV—social views, SP—sympathy.
+The PCA scatter plot in Figure 5 (top left) provides a classical projection of the entire feature set for all politicians, serving as a baseline for comparison.
+In contrast, Figures 7a and b show PCA performed on, respectively, (a) only anonymous features not mapped to avatars or spatial coordinates and (b) all features mapped to spatial coordinates or left anonymous.
+These alternative PCA projections illustrate how the interpretability of the spatial arrangement depends critically on which features are included in the dimensionality reduction.
+In both PCA scenarios, the principal components are determined by mixtures of technical variables (e.g., media activity, voting effectiveness, age).
+No single feature fully dominates the spatial arrangement; both principal components represent blends of the anonymous and/or spatial features.
+Crucially, attributes such as promise fulfillment or group affiliation, which are easily interpreted in the avatar-based view, are not visible in the PCA scatter plot geometry unless they are explicitly included in the projection.
+It highlights the interpretability advantage of spatial avatars: only by explicit mapping can the user guarantee that key variables are visually represented.
+Insights and sensitivity analysis:
+Changing the assignment of features (e.g., switching "fulfillment rate" from an avatar trait to a spatial coordinate or PCA) alters both the visual arrangement and interpretability of the visualization.
+Spatial avatars thus enable flexible, user-driven mapping, supporting either interpretability or structural discovery as needed.
+4.2 Medium-scale illustration: Synthetic soft drinks dataset (N=100)
+The synthetic soft drinks dataset comprises 100 fictional beverages generated to simulate market diversity with three main product groups (classic, premium, and extreme).
+Each beverage is described by intuitive sensory attributes (sweetness, fizziness, overall rating, and color intensity) and technical features (price and caffeine).
+The group_numeric variable encodes the product segment, used only for color coding; each sample is also labeled by name.
+Figure 8: Comparison of classical visualization methods for the synthetic soft drinks dataset.
+Top left: PCA scatter plot; Top right: t-SNE projection; Bottom left: parallel coordinates plot; Bottom right: Chernoff faces grid (6$\times$4) for a subset of beverages, colored by group.
+While PCA and t-SNE reveal some group structure, and parallel coordinates enable visualization of feature distributions, Chernoff faces offer a glimpse into sensory profiles of selected products.
+However, none of these methods provides a direct, interpretable synthesis of both product similarity and profile, as does the spatial avatar approach.
+To provide context for the effectiveness of our approach, Figure 8 shows the synthetic drinks dataset visualized using four classical techniques: PCA scatter plot, t-SNE projection, parallel coordinates, and a Chernoff faces grid for selected products.
+PCA and t-SNE projections highlight clear clustering by product type, with the principal components and t-SNE axes separating the main beverage groups.
+However, interpreting which specific features drive the separation or identifying outliers based on concrete sensory traits remains challenging.
+Parallel coordinates enable a detailed comparison of feature values but quickly become cluttered at this scale.
+Chernoff faces communicate some intuitive traits yet are difficult to interpret en masse and do not encode structural relationships between products.
+Variable assignment:
+Intuitive features (sweetness, fizziness, overall rating, color intensity) were mapped to avatar visual traits (e.g., smile, nose length, skin color, hair color).
+Technical features (price, caffeine) were used for spatial coordinates (X and Y axes).
+The remaining attributes—sourness, sugar, citric acid, CO2 pressure, preservatives, and color intensity—were included as anonymous variables for PCA-based analysis.
+This assignment is reflected in the figures below.
+Figure 9: Spatial avatar visualization of the synthetic soft drinks dataset.
+Avatar features encode intuitive sensory attributes: sweetness (skin color), fizziness (nose length), overall rating (smile curvature), and product group (hair color).
+Spatial coordinates reflect technical properties: price (X) and caffeine (Y).
+This visualization enables immediate comparison of product types and detection of unusual profiles, with both group structure and individual differences readily interpretable.
+PCA loadings and spatial interpretation:
+Principal component analysis was performed under two scenarios: (top) only on anonymous features (not mapped to avatars or spatial coordinates), and (bottom) on all features assigned to spatial coordinates or left anonymous.
+Figure 10 presents heatmaps of PCA loadings and corresponding scatter plots for both cases.
+Figure 10: Effect of variable assignment on principal component analysis for the synthetic soft drinks dataset.
+(top) PCA on anonymous features only.
+(bottom) PCA on all features assigned to spatial coordinates or left anonymous (all except avatar features).
+In both cases, intuitive variables such as overall rating are not directly represented in the geometry unless they are included in the feature set, and their effect may only be visible through color.
+Variable abbreviations: SOU—sourness, SUG—sugar, CIT—citric acid, CO2—CO2 pressure, PRE—preservatives, COL—color intensity, PRC—price, CAF—caffeine.
+Notably, in both PCA scenarios, clear clusters emerge corresponding to beverage groups, as principal components are determined by mixtures of compositional variables (e.g., sourness, sugar, citric acid, preservatives).
+However, while structural groupings are evident, the interpretation of which sensory or technical attributes drive the group separation and direct identification of outliers remains less intuitive than in the spatial avatar visualization (Fig.
+9), where both group membership and profile details are visible at a glance.
+Insights and sensitivity analysis:
+This example demonstrates that, although dimensionality reduction techniques like PCA and t-SNE can recover the underlying group structure when relevant features dominate the variance, only the spatial avatar approach allows for the immediate interpretation of which specific sensory traits differentiate products, providing direct, profile-based insight.
+Switching a feature, such as ”sweetness” or ”fizziness”, from an avatar trait to an anonymous (PCA) feature alters the PCA scatter plot and principal component structure but at the expense of immediate interpretability in the avatar view.
+This flexibility allows users to prioritize either structural discovery or interpretability as needed.
+4.3 Large-scale real-world: Wine dataset (N=1599)
+We evaluated the method on the Vinho Verde dataset [5, 6], which contains 1,599 samples of Portuguese red wines, each characterized by 12 physicochemical and sensory attributes measured in laboratory analysis.
+This dataset is widely used as a benchmark for multivariate modeling and classification tasks in oenology.
+Figure 11: Classical projection methods for the wine dataset.
+Left: PCA scatter plot colored by quality rating; Right: t-SNE projection colored by quality.
+For large datasets such as this, other classical methods (e.g., parallel coordinates or glyph-based Chernoff faces) rapidly become uninterpretable due to severe overplotting and visual clutter and are therefore omitted.
+Classical visualization techniques such as PCA and t-SNE scatter plots (Figure 11) provide a global view of data structure.
+However, for the wine dataset, no distinct clusters or quality-based groupings are apparent; samples of differing quality are dispersed throughout both projections.
+Other approaches, such as parallel coordinates or glyph-based Chernoff faces, are not shown, as they become uninformative for $N\gg 100$ due to overplotting and loss of interpretability.
+Variable assignment:
+Quality, alcohol content, and residual sugar were assigned as avatar features (e.g., smile, face color, nose length), as these are familiar and interpretable for non-experts.
+Four technical variables (volatile acidity, citric acid, residual sugar, sulfates) were assigned to spatial coordinates (X, Y, Z, T).
+All remaining technical features were omitted from the spatial mapping in this scenario.
+Figure 12: Spatial avatar visualization of the Vinho Verde wine dataset.
+Avatar features encode quality, alcohol content, and residual sugar; spatial coordinates reflect volatile acidity, citric acid, residual sugar, and sulfates.
+This mapping enables immediate identification of wines with unusual or extreme property profiles, as well as direct interpretation of both spatial position and visual traits.
+PCA loadings and spatial interpretation:
+For comparison, principal component analysis was performed on all technical features except those assigned to avatar traits.
+Figure 13 presents a heatmap of PCA loadings and the corresponding scatter plot, with samples colored by quality.
+Figure 13: Principal component analysis for the wine dataset based on all technical variables except avatar-assigned ones.
+The first two principal components are mixtures of several physicochemical properties (PC1: volatile acidity 0.57, density 0.40, sulfates 0.44; PC2: residual sugar 0.59, citric acid 0.52, density 0.45).
+No clear clusters or quality-based groupings emerge—high- and low-quality wines are scattered across the projection.
+It highlights the limitation of PCA-based projections for interpretability compared to explicit, user-driven assignments in spatial avatars.
+Variable abbreviations: VAC—volatile acidity, CAC—citric acid, CLO—chlorides, FSD—free sulfur dioxide, SUL—sulfates, DEN—density, SUG—residual sugar.
+In the PCA projection, the principal components reflect combinations of technical features, and the spatial arrangement does not reveal distinct clusters or gradients by quality.
+By contrast, the spatial avatar visualization (Fig.
+12) ensures that critical consumer-facing attributes remain visible and interpretable due to explicit user assignment of axes.
+Insights and sensitivity analysis:
+Assigning ”quality” or ”alcohol” as a spatial or PCA feature, rather than as an avatar trait, would affect the composition of principal components but reduce the immediate interpretability of these key attributes—demonstrating the practical benefit of flexible, user-driven assignment in the proposed approach.
+5 Discussion
+This study introduces a flexible and interpretable framework for multidimensional data visualization, placing the user at the center of the exploratory process.
+The core innovation lies in the semantic division of variables: Users can assign the most meaningful interpretable features (such as quality, popularity, or sensory attributes) to avatar characteristics, while more abstract or technical variables (e.g., chemical or behavioral parameters) are mapped to spatial coordinates.
+Unlike traditional approaches, such as PCA or parallel coordinates—which treat all variables equally and often obscure critical features—our method allows for user-driven, context-sensitive mapping that supports practical interpretability.
+Spatial avatars further extend classical glyph-based visualization by embedding expressive, user-configurable avatars in a fully interactive 3D environment.
+Users can dynamically explore the data cloud, rotate and zoom into clusters, and rapidly test hypotheses by reassigning features to avatars or axes.
+This interactive capacity enables both the detection of structural patterns and the immediate comparison of individual profiles—capabilities not readily available in static projection-based visualizations.
+Empirical results across synthetic and real-world datasets—including fictional politicians, synthetic soft drinks, and the Vinho Verde wine collection—demonstrate that spatial avatars make group structure, outliers, and archetypes directly observable and interpretable.
+In contrast, classical PCA-based projections often mix multiple technical variables into principal components, rendering key features invisible or ambiguous from the user’s perspective.
+By empowering users to control the mapping of features and explore data interactively, the spatial avatar framework bridges the gap between statistical rigor and practical, human-centered interpretation.
+The system, implemented as an open-source plugin for the dpVision platform, is adaptable to various analytical goals, data types, and user preferences.
+5.1 Limitations
+While the proposed approach offers significant advantages, several limitations should be acknowledged.
+The semantic assignment of variables to ”intuitive” or ”technical” categories relies on user expertise and task context, introducing some subjectivity; different users or domains may lead to different choices and, thus, different interpretations.
+The number of avatar visual features is limited, restricting the range of intuitive variables that can be displayed directly, which may necessitate dimensionality reduction or the omission of some information.
+When many variables are treated as ”anonymous” and incorporated via PCA, interpretability can be lost since principal components represent linear combinations rather than original features.
+Moreover, expressive avatars may introduce emotional or cognitive biases, and for very large or high-dimensional datasets, visual clutter or computational limits could arise.
+Finally, variable assignment is currently manual; integrating automated or user-guided assignment tools remains an important direction for further work.
+Although a formal user study has not yet been conducted, the method has been informally evaluated among collaborating researchers.
+These preliminary trials indicated that certain avatar features—such as face color or hair color—are readily distinguished and effective for communicating differences, while others, including smile curvature or subtle facial traits, may be less easily perceived.
+This observation highlights the importance of meticulous glyph design and perceptual testing to enhance interpretability.
+Future work will involve systematic user studies to quantitatively assess the recognizability of individual avatar features and guide further refinements of the visual encoding.
+Addressing these limitations is a natural path for development, including more robust assignment strategies, alternative glyphs to reduce bias, and enhanced support for complex or large-scale datasets.
+5.2 Advantages and utility of spatial avatars
+Compared to conventional dimensionality reduction methods, the spatial avatar approach provides direct, user-controlled interpretability by allowing the explicit mapping of intuitive features to avatar traits.
+It supports immediate visual recognition of unusual or exemplary cases and leverages innate human pattern recognition skills.
+Technical variables projected onto spatial coordinates reveal underlying structures, clusters, or outliers, while avatars simultaneously communicate key user-relevant attributes.
+Interactive mapping and dynamic exploration enable users to iteratively test hypotheses, adjust variable assignments, and visualize the impact on the data structure.
+Empirical case studies (see Section Results) show that spatial avatars reveal patterns and support interpretations that are less accessible or even hidden in standard visualization techniques such as PCA.
+Sensitivity analyses demonstrate how variable assignment impacts interpretability and cluster formation, underscoring the method’s flexibility and value across diverse analytical contexts.
+5.3 Future work
+Several directions for further development are envisaged.
+One approach is to expand the repertoire of avatar types: while Chernoff-like faces take advantage of human abilities to perceive facial traits, alternative glyphs—such as neutral geometric shapes or biologically inspired forms—could reduce emotional bias or adapt to new domains.
+Another area is the automation of variable assignment; future versions could integrate heuristics or machine learning to suggest optimal mapping based on data properties or user priorities.
+Scalability remains an open challenge: dynamic filtering, clustering, sampling, or zoomable detail views may help maintain interpretability in larger data collections, possibly in combination with statistical overviews.
+Ultimately, formal usability studies are necessary to validate the approach in controlled settings and to inform refinements for diverse application areas.
 6 Conclusion
-In this paper, we presented a novel method for visualizing high-dimensional data by combining two natural strengths of the human visual system: the understanding of 3D space and the recognition of facial traits.
-Our approach introduces an important distinction between intuitive and abstract data dimensions and maps them differently for optimal interpretability: intuitive features are assigned to avatar characteristics, while abstract features are reduced and projected into 3D space.
-We implemented the method as a plugin for the open-source dpVision platform and demonstrated its effectiveness on synthetic and real-world datasets.
-Preliminary results suggest that this combination of avatar-based glyphs and interactive 3D navigation offers an intuitive and flexible way to explore complex data structures, particularly when users are not domain experts in mathematics.
-Further development directions are outlined in the next section and include formal usability studies, exploration of alternative glyph styles to mitigate emotional bias, automated feature mapping, and scalability improvements to handle larger datasets.
-By distinguishing between different types of dimensions and tailoring their visual representation, our plugin proposes a human-centered paradigm that could benefit many fields where complex data must be understood and interpreted visually.
-10 FUTURE WORK
-There are several directions in which this work can be further developed.
-One natural extension is to allow for different types of avatars.
-While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation.
-Therefore, alternative glyphs—such as neutral geometric shapes or biologically inspired forms like trees—could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user.
-Such alternatives could reduce the emotional bias associated with human-like faces and allow more neutral data interpretation depending on the application domain.
-Another area of development is the automation of dimension assignment.
-At present, the mapping of data dimensions to either spatial coordinates or avatar features is performed manually.
-In future versions, this process could be assisted or fully automated using heuristics or machine learning methods that take into account feature importance, variance, or user-defined priorities.
-A further line of improvement concerns scalability.
-While the current implementation is effective for moderately sized datasets (typically up to a few thousand points), its perceptual clarity decreases as the number of avatars grows.
-Glyph-based methods such as ours are not ideally suited for very large-scale datasets, as visual clutter and cognitive overload quickly become limiting factors.
-Future versions could integrate mechanisms such as dynamic filtering, clustering, sampling, or zoomable detail views to maintain interpretability in larger data collections.
-Hybrid solutions that combine glyphs with statistical overviews may also be explored.
-Finally, a formal evaluation of the system is planned.
-Although the visualization method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains.
-These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
-11 SUPPLEMENTARY MATERIAL
-The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view.
-The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.
+In this paper, we presented a novel method for visualizing high-dimensional data by integrating two natural strengths of human perception: spatial understanding and intuitive recognition of visual patterns.
+Our approach distinguishes between intuitive and abstract data dimensions, mapping them in different ways for optimal interpretability.
+Intuitive features are assigned to avatar characteristics, while technical features are reduced and projected into 3D space.
+Implemented as a plugin for the open-source dpVision platform, the method was demonstrated on synthetic and real-world datasets.
+The combination of avatar-based glyphs and interactive 3D navigation provides an intuitive and flexible way to explore complex data structures, facilitating the discovery of patterns and individualized interpretation.
+Future work will focus on formal usability evaluation, the development of alternative glyph styles, automated variable mapping, and further improvements to scalability.
+By enabling users to tailor visual representation to their needs, our framework advances a human-centered paradigm for the visual analysis of complex multidimensional data.
+Author Contributions
+Conceptualization, L.L.; methodology, L.L.; software, L.L.
+and D.P.; validation, L.L.
+and D.P.; formal analysis, L.L.; investigation, D.P.; resources, L.L.; data curation, D.P.; writing—original draft preparation, L.L.
+and D.P.; writing—review and editing, L.L.
+and D.P.; visualization, D.P.; supervision, L.L.; project administration, D.P.
+All authors have read and agreed to the published version of the manuscript.
+Funding
+This research received no external funding.
+Data Availability Statement
+The source code for the plugin is available at https://github.com/iitis/N-Dim-view.
+All synthetic datasets used in this study, together with the additional code required to reproduce the experiments and visualizations, are included in this repository.
+The real-world wine dataset used is publicly available and formally cited in the References.
+The source code for the dpVision framework is available at https://github.com/pojdulos/dpVision.
+Conflicts of Interest
+The authors declare no conflict of interest.
+Acknowledgments
+The authors would like to thank Krzysztof Domino, PhD, for his valuable advice and motivation, which supported the creation of this article.
+List of Abbreviations
+PCA Principal Component Analysis
+PC Principal Component
+t-SNE t-distributed Stochastic Neighbor Embedding
+dpVision Open-source Data Visualization Platform
+GUI Graphical User Interface
+SO(2) Special Orthogonal Group in 2D (if used in text)
+SO(4) Special Orthogonal Group in 4D (if used in text)
+SVD Singular Value Decomposition (if used in text)